OpenStudy (anonymous):

A ball is thrown vertically upward with an initial speed of 80 ft/s. Its height after t seconds is given by h=80t-16t^2. a. How high does the ball go? b. When does the ball hit the ground?

OpenStudy (anonymous):

1. \[h = 80t - 16t^2\] Evaluate t from here..

OpenStudy (anonymous):

How? @waterineyes

OpenStudy (anonymous):

I am also not getting this..

OpenStudy (anonymous):

max height is at the vertex put \(x=-\frac{b}{2a}=-\frac{80}{2\times -16}=\frac{5}{2}\) and see what you get for \(h\)

OpenStudy (anonymous):

Wait, sorry I'm still a bit confused. @satellite73

OpenStudy (anonymous):

satellites way is the only way unless you know calculus

OpenStudy (anonymous):

the max height of the ball is at the vertex of the parabola

OpenStudy (anonymous):

Oh, that's probably why then. I haven't learned calculus yet.

OpenStudy (anonymous):